# Dusit Niyato, , Bo Ai, , Khaled B. Letaief

### v1

@@ III-B3 Resource Allocation
Figure 3: Generative AI agent assisted EDoF maximization for rectangular UPA-based XL-MIMO systems with various shapes. We have $L=12\lambda$ and transmitting distance $d=10\lambda$.
IV Case Studies
@@
IV-A Capacity Maximization for Non-Parallel Transceiver
@@
Most studies on XL-MIMO systems focus on an ideal parallel transceiver scenario [15]. However, it is vital to explore the performance for the non-parallel transceiver scenario with a particular rotation angle $\theta$ considered, as varying $\theta$ can significantly impact performance. We investigate how $\theta$ affects capacity in UPA-based XL-MIMO systems, aiming to determine when the system can achieve maximized capacity with different values $\theta$. This analysis is carried out based on the EDoF concept [4], which approximates dominant sub-channels’ count and is closely linked to system capacity.
@@
Applying the generative AI agent framework and the extended database, i.e. a subset of papers from IEEE Xplore and arXiv, we formulate the capacity optimization problem for XL-MIMO systems considering rotation angles, as shown in Fig. 2 (a). We extract the core generated contents by the agent, presented in the figures. The agent can extract important optimization constraints, such as the angle constraint and the transmitting distance constraint. Then, the agent is capable of aiding in the formulation of the capacity maximization problem. Emphasizing the transmitting distance constraint, it is vital for avoiding impractical scenarios and confounding results in the analysis. With the capacity maximization problem defined, we implement the global search method to investigate how $\theta$ affects capacity performance. Simulations reveal that $\theta=0^{\circ}$ can achieve the maximized capacity, indicating parallel transceiver configurations yield the best capacity performance.
@@
Without generative AI agent’s assistance, as seen in Fig. 2 (b), formulating the optimization problem risks missing key constraints, such as the transmitting distance constraint. If the distance is smaller than $L\sin\left(\theta\right)/2$, the receiving plane would touch the transmitting plane, leading to oscillating capacity performance—an impractical and ambiguous result. The employment of the agent is crucial for systematically uncovering modeling ideas and potential constraints, ensuring a more thorough and accurate problem formulation. Notably, generative AI agent not only mitigates the risk of modeling oversight but also enriches the research process, ensuring a broader and more accurate problem-solving framework. With the aid of the agent, in this case study, we explore effects of the non-parallel transceiver on system capacity performance and observe that the parallel transceiver scenario can achieve the maximized capacity performance.
IV-B EDoF Maximization for Various Shapes of Transceiver
@@
We next study the effect of different rectangular shapes of the transceiver on the system performance and provide indispensable insights for practical implementation. Focusing on rectangular plane-based XL-MIMO, we investigate how different rectangular shapes affect EDoF performance, aiming to identify which shape yields the best EDoF performance. The variety of configurations and parameters, however, presents significant challenges for new researchers to effectively understand the underlying motivations and formulate viable problems.
@@
To tackle this challenge, the agent assists the researchers, especially the new researchers, to model and formulate the optimization problem. As illustrated in Fig. 3, by inputting system characteristics and specifying the problems of interest, the agent can adeptly provide the researcher with insightful modeling ideas and comprehensive problem formulation steps. For instance, it suggests simulating various rectangular shapes with constant parameters including antenna number and plane area. With these guidelines, the researchers can systematically explore different rectangular configurations by adjusting the shape ratio. With these guidelines, researchers can be motivated to explore different rectangular configurations by adjusting the shape ratio $0<\alpha\leqslant 1$, uniformly deploying antennas, and calculating EDoF performance for each shape. Simulation results indicate that a square shape ($\alpha=1$) achieves the maximized EDoF performance. This intuitive result can provide vital insights for the practical configuration implementation of next-generation MIMO systems. As demonstrated above, generative AI agent can efficiently facilitate the system modeling and problem formulation, by offering a wealth of insightful ideas and detailed modeling steps, enhancing the efficiency and depth of the research process.
V Future directions
@@
$\bullet$ Explainable AI (XAI): Enhancing the explainability of LLMs is critical for the development of secure, generative AI Agent-assisted wireless networks. This enhancement not only strengthens security and trustworthiness by delineating the AI’s decision-making pathways but also significantly optimizes network functionalities. Specifically, by elucidating the LLMs’ reasoning processes, operators can fine-tune resource allocation algorithms, ensuring that the most impactful factors are prioritized for network throughput and efficiency. For instance, in optimizing XL-MIMO configurations, a transparent understanding of how LLMs evaluate various antenna arrangements or signal processing techniques allows for more informed and effective decision-making. This level of detail in the models’ operational logic enables targeted interventions, directly translating into enhanced network performance and reliability.
@@
VI Conclusions
In this paper, we studied generative AI agent-enabled next-generation MIMO analysis and design. Firstly, the development, fundamentals, and challenges of the next-generation MIMO were reviewed. Then, we investigated generative AI agent and summarized its features and advantages. Next, generative AI agent enabled next-generation MIMO analysis and design were motivated from performance analysis, signal processing, and resource allocation. To clearly showcase the advantages of generative AI agent, we provided two case studies to show how generative AI agent facilitates the next-generation MIMO analysis and design. Finally, some future directions were listed.

### v2

@@ Introduction
Fig. 4: Generative AI agent assisted EDoF maximization for rectangular UPA-based XL-MIMO systems with various shapes. We have $L=12\lambda$ and transmitting distance $d=10\lambda$.
Case Studies
@@
Capacity Maximization for Non-Parallel Transceiver
@@
Most studies on XL-MIMO systems consider an overidealistic parallel transceiver scenario [20]. However, it is vital to explore the performance for the non-parallel transceiver scenario with a particular rotation angle $\theta$ considered, as varying $\theta$ can significantly impact performance. We investigate how $\theta$ affects capacity in Uniform Plane Array (UPA)-based XL-MIMO systems, aiming to determine when the system can achieve maximized capacity with different values $\theta$. This analysis is carried out based on the EDoF concept [21], which approximates dominant sub-channels’ count and is closely linked to system capacity. The EDoF can be computed as $\varepsilon\triangleq{\mathrm{tr}^{2}\left(\mathbf{R}\right)}/{\left\|\mathbf{R}\right\|_{\mathrm{F}}^{2}}$, where $\mathbf{R}=\mathbf{H}^{H}\mathbf{H}$ is the channel correlation matrix with $\mathbf{H}$ being the channel matrix between the transceiver. Then, the capacity can be computed as $C=\mathrm{EDoF}\cdot\log_{2}(1+\frac{\alpha P}{\mathrm{EDoF}^{2}N_{0}})$, where $P$ is the transmitting power, $N_{0}$ is the noise power, and $\alpha$ is the overall channel gain [21].
@@
Applying the generative AI agent framework and the extended database, i.e. a subset of papers from IEEE Xplore and arXiv, we formulate the capacity optimization problem for XL-MIMO systems considering rotation angles, as shown in Fig. 3 (a). We extract the core generated contents by the agent, presented in the figures. The agent can extract important optimization constraints, such as the angle constraint and the transmitting distance constraint. Then, the agent is capable of aiding in the formulation of the capacity maximization problem. Emphasizing the transmitting distance constraint, it is vital for avoiding impractical scenarios and confounding results in the analysis. With the capacity maximization problem formulated, we then apply the global search method to explore how $\theta$ affects capacity performance. Simulations show that $\theta=0^{\circ}$ can achieve the maximized capacity, which means that the parallel transceiver configuration can achieve the best capacity performance.
@@
Without the assistance from generative AI agent, as seen in Fig. 3 (b), formulating the optimization problem may face the risk of overlooking key constraints, such as the transmitting distance constraint. If the distance is smaller than $L\sin\left(\theta\right)/2$, the receiving plane would touch the transmitting plane, leading to an oscillating capacity phenomenon, which is impractical and misunderstanding. Notably, the assistance from the agent is helpful to comprehensively uncover minor modeling ideas and potential constraints. This help can facilitate a more thorough and accurate problem formulation. Notably, generative AI agent not only mitigates the risk of modeling oversight but also enriches the research process, ensuring a broader and more accurate problem-solving framework. With the aid of the agent, in this case study, we study the effect of the non-parallel transceiver on system capacity performance and observe that the parallel transceiver scenario can achieve the maximized capacity performance.
EDoF Maximization for Various Shapes of Transceiver
@@
We next study the effect of different rectangular shapes of the transceiver on the system performance and provide indispensable insights for practical implementation. Focusing on rectangular plane-based XL-MIMO, we investigate how different rectangular shapes affect EDoF performance, aiming to identify which shape yields the best EDoF performance. The variety of configurations and parameters, however, presents significant challenges for beginners to effectively understand the underlying motivations and formulate viable problems.
@@
To tackle this challenge, the agent assists the researchers, especially the beginners, in modeling and formulating the optimization problem. As shown in Fig. 4, by inputting system characteristics and descriptions of the aimed problems to the agent, the agent can adeptly provide researchers with insightful modeling ideas and comprehensive problem formulation steps. For instance, the agent advises to simulate various rectangular shapes with constant parameters including antenna number and plane area. With these descriptions, researchers can be motivated to explore different rectangular configurations by adjusting the shape ratio $0<\alpha\leqslant 1$, uniformly deploying antennas, and calculating EDoF performance for each shape. Simulation results indicate that a square shape ($\alpha=1$) achieves the maximized EDoF performance. This intuitive result can provide vital insights for the practical configuration implementation of next-generation MIMO systems. In this case, generative AI agent can effectively help sort out the key features and important workflows of the intended research problem. As demonstrated above, generative AI agent can efficiently facilitate the system modeling and problem formulation, by offering a wealth of insightful ideas and detailed modeling steps, enhancing the efficiency and depth of the research process.
Future directions
@@
$\bullet$ Explainable AI (XAI): Enhancing the explainability of LLMs is critical for the development of secure generative AI Agent-assisted wireless networks. This enhancement not only strengthens security and trustworthiness by delineating the AI’s decision-making pathways but also significantly optimizes network functionalities. Specifically, by elucidating the LLMs’ reasoning processes, operators can fine-tune resource allocation algorithms, ensuring that the most impactful factors are prioritized for network throughput and efficiency. For instance, in optimizing XL-MIMO configurations, a transparent understanding of how LLMs evaluate various antenna arrangements or signal processing techniques allows for more informed and effective decision-making. This level of detail in the models’ operational logic enables targeted interventions, directly translating into enhanced network performance and reliability.
@@
Conclusions
In this paper, we studied generative AI agent-enabled next-generation MIMO analysis and design. Firstly, the development, fundamentals, and challenges of the next-generation MIMO were reviewed. Then, we investigated generative AI agent and summarized its features and advantages. Next, generative AI agent-enabled next-generation MIMO analysis and design were motivated from performance analysis, signal processing, and resource allocation. To clearly showcase the advantages of generative AI agent, we provided two case studies to show how generative AI agent facilitates the next-generation MIMO analysis and design. Finally, some future directions were listed.